### v1

# A Lean Dataset for International Math Olympiad: Small Steps towards Writing Math Proofs for Hard Problems

Abstract
Using AI to write formal proofs for mathematical problems is a challenging task that has seen some advancements in recent years. Automated systems such as Lean can verify the correctness of proofs written in formal language, yet writing the proofs in formal language can be challenging for humans as well as automated systems. The miniF2F dataset, the community’s benchmark for automated theorem proving, has 20 IMO problems in its testing set, yet formal proofs are available only for 7 of these problems (3 of which are written by mathematicians). The model with best accuracy can only prove 4 of these 20 IMO problems, from 1950s and 60s, while its training set is a secret. In this work, we write complete, original formal proofs for the remaining 13 IMO problems in Lean along with 3 extra problems from IMO 2022 and IMO 2023. This effort vastly expands the availability of proof currently in the public domain by creating 5,150 lines of lean proof. The goal of the paper is to pave the way for developing AI models that can automatically write the formal proofs for all the IMO problems in miniF2F and beyond. In this pursuit, we devise a method to decompose the proof of these problems into their building blocks, constructing a dataset of about 900 lemmas with 25,500 lines of Lean code. These lemmas are not trivial, yet they are approachable, providing the opportunity to evaluate and diagnose the failures and successes of AI models. We then evaluate the ability of GPT-4 in writing formal proofs for these lemmas. This involves zero shot prompting, Chain-of-Thought reasoning with expert feedback and lemma retrieval. In evaluating the responses, we also analyze the confounding factor of LLM’s ability to write the proofs in natural language vs Lean language. From all these evaluations, two clusters of IMO problems emerge: one group are the problems prior to 1970s where LLMs have a higher success rate in writing a proof for their decomposed lemmas, they can explain a correct proof in natural language, and when they fail, expert feedback and lemma retrieval improves their answers. The second group of problems are the IMO problems from the more recent years, their proofs are longer, subjects are more challenging, the success rate of LLM in proving the decomposed lemmas is small, LLM cannot explain a correct proof for them in natural language, and finally, lemma retrieval and CoT rarely have a positive impact on the accuracy of their formal proofs.
1 Introduction
Machine learning methods have made significant progress in many different domains. This progress can often be described as the automation of learning from data. Availability of high quality data is crucial for the community to develop better models and methods. ImageNet [10], for example, played a significant role in the development of ML methods in computer vision. More recent models such as GPT and Llama rely on training set purportedly almost as large as the entire contents of the Web. On the other hand, the failures of these models on topics such as math are often explained via lack of high-quality materials on those topics in their training set. This issue is partially addressed by curation of relevant data, e.g., [17] recruits human labelers and AlphaGeometry [29] generates synthetic data. For the topic of automated theorem proving in formal language, high quality data is so scarce, to the extent that the formal proofs are not available for almost half of the problems in the miniF2F dataset, the benchmark of the community. So, when automated methods fail to write a correct proof for the problems in the benchmark dataset, it is not clear where they have failed, and what the correct answer looks like. In this work, we study 16 IMO problems, in number theory and algebra. 13 of the 16 problems are from the miniF2F dataset and 12 of the 16 do not have a formal proof available in the public domain. Moreover, for all the 40 IMO problems in the miniF2F dataset, we contribute the complete, original, formal proofs in Lean language.
Solving mathematical problems is a challenging task for machine learning. There are many motivations for doing math with AI as we will discuss further in Appendix B. One motivation is the continuation of creating machines that can outperform humans in a given task, such as in the game of chess [19] and two decades later in the game of Go [25]. Currently, there is a 10 million dollar prize for an AI system that can participate in the International Math Olympiad and win a gold medal: the AIMO Prize [1]. There is also the IMO Grand Challenge [2] which specifically requires the model to write its proof in Lean language. Lean is a functional programming language and a proof assistant [9] used mostly by mathematicians to write and verify mathematical proofs. The Lean system makes it possible to verify the correctness of a proof instantly. This has given rise to a field known as formal mathematics. Lean is also used by Amazon AWS for formal verification of software [8].
1.1 IMO: an unknown testing set that gets harder every year
Challenging an AI model to participate in an IMO and win a gold medal has certain implications, especially regarding the evaluation. First, the problems presented in an IMO are revealed to participants on the day of the exam. Doing well in the IMO requires a mastery of the subjects presented in it. Merely studying the problems from the previous iterations of the Olympiad is not sufficient as the problems tend to become harder year by year. Second, the problems that appear in IMO go through a selection process that ensures some degree of novelty. As a result, these problems are usually not present in text books. It follows that direct retrieval from the textbook solutions is usually not sufficient to prove the new IMO problems. Third, the exam is timed. A participant must complete the exam within 9 hours.
These conditions, in the language of machine learning, implies being evaluated on an unknown testing set with samples that are more challenging than the samples in the training set. In the machine learning literature, however, having standard benchmarks with fixed and accessible testing sets is essential for a large community to be able to work towards common goals. This approach of using a fixed testing set has worked reasonably well in some ML fields [22], mostly because the training sets of the models are fixed, too, and in creating these datasets, specific procedures are followed to ensure samples from training sets do not appear in the testing sets. However, with the rise of large language models, it has become increasingly difficult to evaluate how much of a model’s success is merely direct retrieval of the correct answer from its training set. Although for solving mathematical problems, the miniF2F [34] dataset is the common benchmark with a fixed testing set, the training set of LLMs is sometimes unknown; when it is known, it includes all the relevant content available on the Internet, (e.g., all the available code on Github, arXiv, Stack Exchange, and elsewhere), which is an enormous corpus, hard to search through and compare against the testing sets. When evaluating LLMs developed on such training sets, we often use a fixed testing set. But, IMO evaluates on an unknown testing set that gets harder every year.
To prepare an AI model that can do well in IMO, we need to rely on rigorous evaluation techniques that can tell us whether the model can generalize well on an unknown testing set. In some applications, the distinction between generalization and retrieval might be besides the point. As long as a model performs the task correctly, one might not be too keen to evaluate what percentage of its correct answers is a direct retrieval from its training set. However, when we are developing a model that is intended to be evaluated on an unknown testing set, we will need to rely on clearer evaluation methods.
1.2 Where we are
Let us now take a closer look at the current benchmark: miniF2F dataset. This dataset consists of mathematical theorems with varying degrees of difficulty. Out of the 244 theorems in its testing set, 20 of them are IMO problems, and the difficulty levels of those IMO problems also vary since they are taken from years 1959 to 2019. In terms of complexity, some of the theorems in the dataset can be proved with one line of formal proof while others may need more than 500 lines of formal proof (see Table 1). Success in solving a high percentage of the less complex problems is not the same as success in solving the same percentage of the more complex problems. Hence, merely reporting the accuracy by the percentage of the testing set is not insightful enough.
LEGO-Prover [30], the current state of the art method on miniF2F, is able to prove only one of the IMO problems in this dataset: the one from 1959, which is the first problem of the first IMO problem implying that it is the easiest of all IMO problems proved in less than 10 lines of code. Other ML methods have proved more of the IMO problems in the miniF2F, e.g., the HTPS method by [16] claims proving 10 IMO problems (not all from the miniF2F), but those problems are either from 1960s or the ones for more recent years have a simplified or erroneous problem statement. The method by [21] also proves two adapted (simplified) IMO problems.
1.3 Our approach
For the majority of the IMO problems in miniF2F, even a human-written formal proof is not publicly available. In this work, we look into these specific problems along with 3 more challenging problems from IMO 2022 and 2023.
We provide the formal proofs for these problems and further evaluate the ability of GPT-4 in writing these proofs. We decompose each of the formal proof steps into small lemmas and report that GPT-4 cannot write a correct formal proof for almost all of the lemmas. Writing the formal proof for these problems can be viewed from the perspective of planning and execution: one would need to prove a certain chain of lemmas (execution), and put those lemmas together (planning) in a coherent way to compose the final proof. The planning aspect of writing formal proofs is challenging, and in the literature, often the plan, i.e., the proof sketch or the approach, is taken from the informal proofs written by humans [14].
We diagnose the failures of GPT-4 by evaluating and labeling its proofs manually. We further guide the model by pointing out its mistakes. Giving feedback to the model is mostly unfruitful, especially on the topics that GPT-4 seems to be less educated. Finally, we perform source criticism by comparing the proofs written by GPT-4 against the contents of the Web, as a proxy for its training set. This systematic study provides insights as to what fraction of the answers of GPT-4 can be considered direct retrieval from its training set. We report a strong correlation between the correctness of the answers of GPT-4 and the existence of proofs on the Web.
1.4 Our contributions
  1. 1.
    We provide the formal proofs for 12 IMO problems that do not have a formal proof in Lean. Nine of these problems are from the miniF2F dataset. The other three are from IMO 2022 and 2023. We specifically chose the problems that do not have a formal Lean proof in the public domain. Our proofs are written in Lean3 and 4 and consist of 3,100 lines of code.
  2. 2.
    We decompose the proofs for 12 problems into small lemmas, creating a dataset of more than 900 lemmas that are still challenging for AI models, yet they are easier and more approachable. Our method can be extended to all the IMO problems that we have formalized.
  3. 3.
    We evaluate the ability of GPT-4 in writing the formal proofs for the lemmas in on our dataset. We perform a thorough human review of the GPT-4’s informal proofs as well as its formal proofs. For incorrect proofs, we consider Chain-of-Thoughts techniques to guide the GPT-4. We further quantify the human effort required to repair the incorrect proofs.
2 Dataset of approachable lemmas: Stepping stone towards solving challenging IMO problems
Table 1 describes the 12 IMO problems that we study in this paper. The concepts used in these problems are described in further detail in Table C1. In appendices D and F, we provide the datasheet and license information about our dataset.
We report that GPT-4 cannot write a correct formal proof for any of these problems. In natural language, though, it can describe a sensible approach for proving the first 6 IMO problems. For the IMO problems between 1984 and 2023, its approach in natural language is also wrong.
Table 1: IMO problems formalized and studied in this paper
# Year Problem Topic in miniF2F Lean proof available # of lemmas # of lines of Lean4 code
1 1959 p1 number theory Yes No 3 9
2 1960 p2 algebra Yes No 8 40
3 1962 p2 algebra Yes No 14 60
4 1964 p2 algebra Yes No 9 50
5 1965 p2 algebra Yes No 73 210
6 1983 p6 algebra Yes No 55 180
7 1984 p6 number theory Yes No 68 380
8 1992 p1 number theory Yes No 91 480
9 1997 p5 number theory Yes No 122 390
10 2022 p2 algebra No No 61 260
11 2022 p5 number theory No No 266 640
12 2023 p4 number theory No No 137 450
total 907 3,149
Making these formal proofs accessible to the research community can be helpful in different ways. First, every one, including non-mathematicians, will know the work it takes to write a formal proof for these problems. Second, these formal proofs can also be used to identify possible mistakes or shortcomings in the informal proofs that might be circulating. For example, [5] reports that GPT-4 can write a correct informal proof for IMO 2022 P2. However, when we write the formal proof for the same problem, it becomes obvious that the informal proof was incorrect and missing a crucial step. Third, some of the formal proofs in the literature simplify the original IMO problem by altering the problem statement, but we remain faithful to the original problem statements of IMO. For example, [31] showcases the proof for two shortlisted IMO problems. In the first problem, IMO 2009 P3, there is a mistake in the problem statement, and the proof merely uses the error to prove False. For the second problem, IMO 2016 P5, the statement is altered from proving “there exists infinitely many solutions" to “there exists a solution".
2.1 Decomposition of the proofs into lemmas
In the next step, we decompose the proofs for each of these problems into small lemmas. For example, the proof might involve showing that given a set of constraints, natural number $p$ cannot be larger than 4. Then, use another series of deductions to show that $p$ cannot be smaller than 2. And then use these two results to conclude that $p$ must be either $2$ or $3$. Appendix E provides a more detailed discussion about decomposition of lemmas and some upper bounds on how many proofs can be extracted from $n$ lines of proof. Moreover, Table E2 provides some statistics about the length of the proofs for the lemmas in our dataset.
All of the 907 lemmas in our dataset have a formal proof in Lean. Some of the easier lemmas are as the following:
  • $q$ and $r$ are integers and their product is 11, prove that $q$ is either ±$\pm 1$ or ±$\pm 11$
  • $q$ and $r$ are natural numbers and their product is 5, prove that either $q$ or $r$ must be 1.
  • $p$ is a natural number greater than 4, prove that rational number $\frac{p}{(p-1)}\leq 4/3$
  • $k$ is natural number greater than 4, prove that $k<2^{(k-2)}$
Our library of lemmas also includes harder problems such as proving injectivity of functions, points about rational numbers, casting, divisibility. Here are some examples:
  • $x$ and $y$ are positive natural numbers where $y<x$, and we have $(x^{y})^{2}=y^{x}$, prove that $y^{2}$ divides $x$
  • $a$ and $b$ are natural numbers and $p$ is a prime number where $a^{p}=b!+p$ and $p\leq b$, prove that $p$ divides $a$
  • $b$ is a natural number and $p$ is a prime number, where $p\leq b$ and $p^{p}=b!+p$, prove that $p$ must be less than 5
The lemmas that are harder to solve are also broken into smaller pieces, so the dataset includes both the hard version of such lemmas plus the smaller lemmas needed to prove them. Later in section 4, we will go through a case study of such lemma to give a better understanding of the dataset and what it takes to prove the lemmas.
2.2 Excluding the problems that native Lean methods can prove
We intentionally exclude lemmas that can be solved automatically via Lean’s native methods such as hint, linarith, exact?, simp, omega, ring, norm_cast and norm_num. This is because we consider those solvers efficient and we do not find it essential to develop automated methods doing the same job as those solvers. We have empirically seen that GPT-4 fails to write a correct formal proof for many of those easy problems as well. Nevertheless, our main objective here is not to evaluate GPT-4, but to create a dataset that can be used as a stepping stone for automated systems. If a task can already be performed adequately and fast by some open source solver, our view is that it would be better to use such a solver and even incorporate it into a more sophisticated system.
3 GPT-4’s ability to solve the problems
3.1 Prompting GPT-4
Here, we prompt GPT-4 with a statement requesting it to write a formal proof for the problem, instructing it to think step by step, explain the theorem, the approach, and the proof, following prompting techniques suggested by [15]. We also provide the formal Lean statement of the problems in the prompt so that issues about the formalization of the theorem statement will not arise. Furthermore, we crosscheck whether GPT-4 can explain the formal Lean statement in English, and whether the natural language description of the GPT-4 for each lemma was correct (see Figure G1). Therefore, providing the problem statement in Lean does not create confusion for the model as it can describe it correctly in English. Please see Appendix G to see the exact prompt structure and some examples.
Table 2 shows the evaluation of the responses of GPT-4 in natural language (NL) and in Lean. GPT-4 defaults/prefers to write formal proofs in Lean3 rather than Lean4. Hence, we evaluate its formal proofs directly in Lean3. The last column of this table evaluates whether there is a match between the proof written in NL and the one in Lean. GPT-4 may describe a proof in NL, but follow a different approach when writing the proof in Lean. For example, for the three lemmas corresponding to IMO 1959 P1 (the first row in Table 2), GPT-4’s proof in NL is correct for all the lemmas (100% in "Correct proof in NL"). However, GPT-4’s formal proof for one of these lemmas do not match the approach described in NL: 67% in "Match between NL and Lean". Indeed, GPT-4’s formal proofs for all these three lemmas have one or more mistakes: 0% in Correct proof in Lean.
Table 2: Analyzing the GPT-4’s responses to our lemmas
# Problem Correct proof in NL Correct proof in Lean Match between NL and Lean
1 1959-p1 100% 0% 67%
2 1960-p2 25% 25% 75%
3 1962-p2 73% 60% 60%
4 1964-p2 50% 40% 40%
5 1965-p2 36% 21% 15%
6 1983-p6 32% 13% 12%
7 1984-p6 6% 0% 5%
8 1992-p1 0% 0% -
9 1997-p5 0% 0% -
10 2022-p2 0% 0% -
11 2022-p5 0% 0% -
12 2023-p4 0% 0% -
Moreover, Table 3 shows a more detailed evaluation of the formal proofs written by GPT-4. If a proof is completely correct and can pass the Lean system, it will be counted in the column "No error". Otherwise, a proof may have one or more of the following issues: hallucinations, wrong approach, wrong implementation, and minor errors.
"Hallucination" refers to cases where the proof contains a lemmas or tactic that does not exist in the Mathlib library and it is also not defined separately in the proof. This happens more often when the model is uneducated on a topic. In such cases, GPT-4 usually assumes that there is one or a few magic lemmas in the Mathlib library that can prove the given lemma outright or in a few steps.
"Wrong approach" refers to cases where the approach to prove a lemma is wrong. This means that either some of the individual proof steps are wrong or the steps do not logically lead to what the model wants to prove.
"Wrong implementation" refers to cases where the approach itself is correct but the model does not use the correct procedures or it does not call the correct lemmas from the Mathlib library.
Finally, "minor errors" refers to cases where the model uses a correct lemma or tactic from the Mathlib library but it does not provide the correct arguments to it, or it misses some of the arguments.
A given proof may have several lines of code, and many of these issues may arise. For example, a proof may have a few hallucinations, a wrong implementation, and also a few minor mistakes.
Table 3: Analyzing and grading the errors in the Lean proofs written by GPT-4
# Problem No error One or more types of error
Hallucinations Wrong approach Wrong implementation Minor errors
1 1959-p1 0% 67% 33% 67% 67%
2 1960-p2 12% 62% 75% 88% 88%
3 1962-p2 33% 40% 40% 47% 67%
4 1964-p2 10% 40% 50% 60% 90%
5 1965-p2 11% 71% 65% 79% 94%
6 1983-p6 0% 84% 91% 100% 100%
7 1984-p6 0% 100% 100% 100% 100%
8 1992-p1 0% 100% 100% 100% 100%
9 1997-p5 0% 100% 100% 100% 100%
10 2022-p2 0% 100% 100% 100% 100%
11 2022-p5 0% 100% 100% 100% 100%
12 2023-p4 0% 100% 100% 100% 100%
3.2 Interacting with GPT-4: Hinting and Chain of Thoughts
We then try to improve the response of GPT-4 by interacting with it. For each incorrect proof, we provide feedback, pointing out the mistakes, and asking the model to fix the issues. If the approach is wrong, we mention that the approach is wrong. If there are hallucinations, we point out such mistakes as well. The results are presented in Table 4.
Table 4: Interacting with and hinting GPT-4 to improve its proof
# problem % correct zero shot % correct after 5 rounds of feedback % correct after 10 rounds of feedback
1 1959-p1 0% 67% 67%
2 1960-p2 12% 25% 25%
3 1962-p2 33% 47% 47%
4 1964-p2 10% 30% 30%
5 1965-p2 11% 17% 17%
6 1983-p6 0% 7% 8%
7 1984-p6 0% 8% 8%
8 1992-p1 0% 0% 0%
9 1997-p5 0% 0% 0%
10 2022-p2 0% 1% 1%
11 2022-p5 0% 0% 0%
12 2023-p4 0% 0% 0%
We interact with GPT-4 at most 10 times. If the model does not succeed in writing a correct proof after 10 rounds of feedback, we stop the process. There are papers in the literature that go as far as hinting the GPT 100 times. We find such an approach labor-intensive and not scalable. It is worth noting that 10 times can already be costly without bringing fruits. Moreover, we have observed that within that range of 10 times, GPT-4 has made no significant improvements.
Later in this paper, in the section on source criticism, we show that most of the GPT-4’s correct answers are already available on the Internet. Hence, a search engine that can retrieve from the contents of the Mathlib library or simply the contents on the Internet may be more efficient than interacting with GPT-4 many times over.
The results further suggest that GPT-4 can be considered uneducated on many of the subjects listed in Table C1, especially the topics related to IMO problems between 1984 to 2023. One might imagine that having LLMs trained on these topics may have the potential to address some of the shortcomings. However, since the search space and the number of combinations of lemmas in topics such as number theory are immensely large, unlike the problems in geometry, it remains a questions whether the best model to write formal proofs for problems in number theory and algebra will be an LLM or some other method that primarily relies on searching the existing lemmas in the Mathlib library.
We also perform manual proof repair as described in Appendix I.
4 Case study: Insights about the dataset
To provide a better understanding of the dataset and GPT-4’s abilities, in this section, we provide a case study on a small portion of the proof from the IMO 1997 P5. Note that in this section, we follow a chain of simplifications of the problem for the purpose of demonstration. Unlike this section, the results in section 3.2 follow a specific protocol of asking GPT-4 to prove one specific lemma, and repeating the same request while providing feedback about the mistakes of the previous response.
Out of the 390 lines of Lean proof for this problem, we zoom in on only 80 lines of the proof. Here is the lemma:
[…]
In the first attempt, GPT-4 gives a very long answer, reviewing all the items in the lemma, but does not provide any proof, not in natural language and not in Lean. It does not know how to approach the problem.
For the proof, one needs to first show that $2y^{2}<x$. After proving this inequality, we can rely on prime factorization of both sides of the inequality to prove the target of the lemma. To guide the GPT-4, we directly ask if it can prove $2y^{2}<x$. Again, it goes on reviewing the statements line by line, providing a large amount of explanations, but, eventually, does not find a way to prove it – again no proof in Lean and no proof in natural language.
We then go back to the first statement and take it for granted that $2y^{2}<x$. This time, GPT-4 attempts to provide a proof, a simple proof relying on a lemma called nat.dvd_of_factorization_eq. Such lemma does not exist in the Mathlib library so its use can be considered a hallucination, perhaps out of desperation. We point out this issue to GPT-4, which leads it to write another proof including this:
rw nat.factorization_mul_of_coprime (nat.coprime.symm (nat.prime.coprime_pow_two_of_ne (ne_of_lt hyx) (nat.prime_of_mem_factorization (exists_mem_of_ne_empty (factorization_pos_of_pos y_pos)))) x_pos) at $h_{2}$,
This line contains many non-existent lemmas. We ask GPT-4 to explain the lemmas one by one which it does happily not realizing that they are hallucinations. As an example, here is the description of GPT-4 for the non-existent lemma nat.prime.coprime_pow_two_of_ne:
"This lemma asserts that if $a\neq b$, and a is a prime, then $a$ is co-prime to $b^{2}$. This is based on the property that a prime number does not divide any power of another distinct number."
GPT-4’s explanation is completely wrong and implies a lack of understanding about the topic.
We further simplify the problem by granting the statement that
$x\bullet 2\bullet nat.factorization\;y\leq x\bullet nat.factorization\;x.$
This statement by itself needs several lines of proof, but we grant it to GPT-4 to make the problem easier. GPT-4 again goes back to its original hallucination of the lemma nat.dvd_of_factorization_eq. We inquire about this lemma and GPT-4 explains that if two numbers have equal prime factorizations, then those numbers will divide each other. We ask GPT-4 if it has been able to establish that $y^{2}$ and $x$ are equal. It answers no and apologizes. Then we point out that nat.dvd_of_factorization_eq does not even exist. It apologizes again.
Finally, to step aside from prime factorization, we simplify the problem further by adding a new statement $(y^{2})^{x}|x^{x}$, and ask GPT-4 to prove $y^{2}|x$. This time, it ignores the new statement, provides a wrong proof in natural language, and does not write any proof in Lean.
What we can conclude from this case study is that GPT-4 does not have adequate knowledge about this topic, neither in natural language, nor in Lean. Anytime it tries to come up with an answer, it hallucinates lemmas that can prove the problem outright, but those lemmas do not exist, and at the same time, they are incorrect or inapplicable. So, it is not viable to prove those lemmas and then use the rest of the GPT-4’s proof.
This case study is insightful, not as a failure case of GPT-4, but as a showcase of the gap between the GPT-4’s abilities and what it takes to write the formal proof for the IMO 1997 P5. This lemma was just a small portion of the proof for this IMO problem, and until we have a model that can prove this lemma, we will not be able to write the complete proof for this IMO problem.
5 Related work and a discussion on avoiding training/testing set contamination for better evaluation of model generalizations
One way to address the evaluation of generalization abilities of an AI model is to test them on challenging problems where the correct answer is not publicly available. For example, FunSearch [23] works on combinatorial problems such as bin packing and comes up with a heuristic algorithm that is better than the best algorithm available in the literature. The last improvement on this problem was made decades ago, so it is clear that this algorithm was not copied from somewhere else.
AlphaGeometry [29], on the other hand, evaluates its performance on all IMO problems in geometry from 2000 to 2022. To avoid training/testing set contamination, it trains its language model on a synthetic dataset generated automatically using an SMT solver. This method of using synthetic data addresses many of the evaluation concerns, but the starting point for creating the synthetic data is geometric figures, and then they exhaustively apply all the possible geometric lemmas to those figures. This approach works for geometric problems, especially when a capable SMT solver is available. Moreover, these geometric problems are decidable unlike the ones we study in this paper. For problems in number theory and algebra, the approach of exhaustively applying all the possible lemmas to a given state is not practical, as the number of applicable lemmas is much larger compared to geometric problems. Moreover, AlphaGeometry considers proving the problems in a system like Lean to be difficult and uses a more simple symbolic system developed by the mathematicians.
LeanDojo [32] observes that: "the common practice of splitting theorems randomly into training/testing has led to an overestimated performance in the previous papers. LLMs can prove seemingly difficult theorems simply by memorizing the proofs of similar theorems during training." To remedy this issue, it proposes a new data set in which the testing set is designed to be challenging compared to the training set.
Outside the domain of formal mathematics, Skill-Mix [33] proposes an evaluation technique in which prompts are designed to include a combination of skills (such as metaphor, red herring, and common knowledge physics). Combining skills in this way aims to prompt the model to come up with a sensible answer that most likely does not already exist in the training set of the model. If the model were to produce a sensible answer, it would need to draw from various parts of its training set. Skill-Mix’s study provides positive evidence that GPT-4 can occasionally provide sensible answers to prompts combining a small number of those skills. But for most prompts, GPT-4 does not succeed.
Other studies use pretrained language models to prove theorems, i.e., LLMs that are trained on the contents of the Web. Some of these papers acknowledge that the answers to some of the testing problems might have leaked into the model training set, for example, [11] reports: "there is the potential for proofs from the test set to have leaked into the LLM pretraining data. While the pretraining data for the Minerva LLM at the base of our models does not include the PISA dataset, it does contain code that may include some Isabelle/HOL proofs found in PISA. This should be kept in mind when interpreting the results."
Some other studies do not try to make a clear separation between the training and testing set; neither do they try to distinguish between the generalization and possible retrievals from the the training set or the prompt. For example, LEGO-prover [30] uses ChatGPT to convert a human written proof into a proof sketch for each of the problems in miniF2F, and then again uses ChatGPT to write the formal proof for each of the steps in the proof sketch. When prompting ChatGPT, it tags the possible relevant lemmas and their proofs in the prompt so that GPT can use them if needed. These additional lemmas are taken from a separate library. For each problem in miniF2F, LEGO-prover performs up to 100 proof attempts with ChatGPT costing around 300 dollars. At the end of the day, it is not clear how many of those success rates are randomly exhausting all possible combination of lemmas and how many of them might be direct retrieval from the contents of the prompts or contents of the ChatGPT’s training set. Moreover, since ChatGPT is a model that evolves over time and it also is not open source, concerns about the reproducibility of the results arise.
6 Conclusion and Future Work
In this work, we presented a dataset of more than 900 lemmas as a stepping stone towards writing automated formal proofs for challenging IMO problems in number theory and algebra. We devised a systematic method to create a large data set of approachable lemmas formalized in Lean. These lemmas are the building blocks of proofs for IMO problems. They are considerably less challenging than IMO problems, yet GPT-4 struggles to prove almost all of them. Moreover, we performed an extensive human evaluation of GPT-4 responses to diagnose the issues this model encounters when it tries to write a Lean proof for these problems. Our experiments on these lemmas allowed us to diagnose and evaluate the obstacles towards proving the unsolved IMO problems in the miniF2F dataset, and beyond, including the more challenging ones from 2022 and 2023. Developing models that can prove the lemmas in these datasets would brings us closer to models that can participate and do well in IMO.
Appendix A Limitations
This work is a step towards writing formal Lean proofs for challenging problems in mathematics, specifically IMO problems. The IMO problems we chose cover some of subjects covered in International Math Olympiad, mostly the ones related to number theory and algebra. Topics such as combinatorics and geometry are not covered here. Moreover, our dataset does not cover all the topics under the category of number theory and algebra, though, we consider our work a major step in that pursuit.
Writing a proof for an IMO problem on number theory and algebra requires both planning and execution of the steps in the plan. One has to look both at the big picture and also at the building blocks that build on top of each other to create the proof. In this paper, our focus was largely on the building blocks and small lemmas, but to have an automated system that can prove such IMO problems one would need to give a considerable attention on planning a viable proof sketch and then putting the building blocks together.
Our dataset is in Lean. It may be converted to other languages such as Isabelle as well. We are familiar with Isabelle and we may perform this conversion if we see interest from the AI researchers who primarily work with Isabelle.
We do not foresee any negative societal impact specifically resulting from this work. Automation sometimes eliminates or reduces jobs. At the same time, it creates new jobs. Developing models and methods that can solve mathematical problems may have such effects on the labor market, but this is not necessarily a negative impact.
Appendix B Other motivations
Earlier, we outlined the motivation for developing AI models that can perform well in IMO. We have noted some of the other immediate and practical motivations for building such systems. Here, we elaborate on what has been delineated above.
We have mentioned Amazon’s use of Lean system for formal verification of software. Besides Amazon, Lean is being used by a rather large community of mathematicians to write formal proofs for a variety of mathematical problems. For example, Terry Tao and a team of other mathematicians are currently formalizing the prime number theorem (i.e., the asymptotic distribution of the prime numbers among the positive integers) in Lean. For an ordinary theorem, it is reported that writing a formal proof in Lean may take up time more than 10 times longer than writing the same proof with pen and paper. For the more complicated topics such as the prime number theorem, it may take months to build the necessary background in Lean language to prove the ultimate theorem. However, the correctness of a proof written in Lean can be verified instantly, but the proof with pen and paper may have shortcomings or even mistakes that are hard to detect. For complicated problems, finding a human reviewer to verify a proof written with pen and paper may not be easy.
As the Mathlib library grows, and as more lemmas become available in the library, lemmas that can be utilized in writing new proofs, the amount of time for writing a proof for simple problems is expected to reduce. In writing these proofs, mathematicians use their knowledge of Lean and Mathlib to write the proofs. Machines may also be helpful in reducing the time for the formalization of proofs [6]. GPT-4 has not appeared as a useful tool for writing formal proofs for novel mathematical problems. But, there is a desire and demand for automated systems that are capable to help the mathematicians in writing the formal proofs. In Lean, there are native methods that can suggest tactics or even prove some theorems by applying existing lemmas. Researchers also developed LLMs that can suggest tactics aiming to help mathematicians at writing the formal proofs, e.g., the Copilot by [26]. However, such LLMs appear to be in their infancy, e.g., Copilot only proves easy lemmas that are designed for new learners of Lean.
On the other hand, for mathematicians, there are always harder theorems and conjectures that can be taken up to be formalized. In other words, there is no limit on how large a mathematical library can grow. Most recently, Kevin Buzzard started a project to formalize the proof for Fermat’s last theorem in Lean, and this project is expected to run for a few years [7].
The process of peer review in these fields of mathematics is largely concerned with verification of correctness of proofs. So, when a new paper is written and its proofs are formalized in Lean, the review process becomes much easier and faster. The large memory and computational power of computers can also be beneficial in proving novel theorems that require vast exploration of possibilities and hard for a human to prove and verify.
In summary, having automated systems that can follow a certain logic, utilize a library of existing lemmas, and prove novel mathematical theorems would have a significant impact on mathematical research and possibly other fields such as design of algorithms, formal verification, etc.
Appendix C Mathematical topics used in our dataset
Table C1 shows the topics used in each of the IMO problems we have studied in this paper. For an automated system to be able to prove the lemmas in our dataset, it will need to utilize these topics and concepts.
Table C1: Topics utilized in the IMO problems formalized and studied in this paper
# Problem Types Concepts used in the proof
1 1959-P1 Natural divisibility
2 1960-P2 Real algebra, inequalities
3 1962-P2 Real, Rational quadratic discriminant and roots, quadratic factorization, algebra, inequalities
4 1964-P2 Real algebra, inequalities
5 1965-P2 Real linear system of equations, algebra, inequalities
6 1983-P6 Real algebra, positive and negative inequalities, bounds
7 1984-P6 Natural power inequalities, divisibility, algebra, prime factorization, greatest common divisor
8 1992-P1 Natural, Integer, Rational factorization, primes, group theory, casting, exponential growth, absolute values
9 1997-P5 Natural, Real prime factorization, divisibility, group theory, rings, logarithm, casting, induction
10 2022-P2 Real algebra, inequalities, order, logic
11 2022-P5 Natural, Integer congruence, primes, lifting the exponent lemma, factorial, finite sets, big operators, divisibility, functions
12 2023-P4 Natural, Real finite sets, functions, big operators, casting, weighted geometric mean
Appendix D License and authors’ responsibility statement
We plan to release our dataset under the MIT license on GitHub and also via Croissant. Authors bear all responsibility in case of violation of rights.
Appendix E Decomposing a proof into its building blocks and extracting new lemmas
In this appendix, we explain the procedures and derive some bounds abound extracting new lemmas from an existing proof.
The lemmas in our dataset were created by breaking down the proofs of the IMO problems. These proofs usually have specific structures that we exploit when breaking them down.
Later, we will consider an extreme case and derive some bounds on the number lemmas that can be extracted from an n line proof. But, before that, we explain, how we exploit the structure of the proofs.
E.1 Breaking down the structure of a proof
Let us consider a theorem with $n$ lines of proof where $n>3$. This proof may contain $k$ intermediate hypotheses that are proved one after another prior to proving the main statement of the theorem. These $k$ hypotheses, each have a proof of their own, and each may have intermediate hypotheses of their own, as well. Here, we only explain one level of breaking down the hierarchy.
Let us use the case study in Section 4 as an example. Given a set of hypotheses (explained on page 7), we want to prove $y^{2}$ divides $x$. First, we can consider each of the intermediate hypotheses and define them as a separate lemma. The intermediate hypotheses in the proof are:
  • Lemma 1: prove $y^{2}<x$
  • Lemma 2: prove $2*y^{2}<x$
  • Lemma 3: prove prime factorization of $(y^{2})^{x}$ <= prime factorization of $x^{x}$
  • Lemma 4: $(y^{2})^{x}$ divides $x^{x}$
Each of these steps are taken as a lemma since their proofs satisfy the two criteria mentioned before (at least 2 lines of proof and not provable by native provers in Lean).
Moreover, new lemmas can be constructed by the addition of the above lemmas to the hypothesis space of the original theorem.
  • Lemma 5: grant lemma 1 to the original problem and prove $(y^{2})^{x}$ divides $x^{x}$
  • Lemma 6: grant lemmas 1,2 to the original problem and prove $(y^{2})^{x}$ divides $x^{x}$
  • Lemma 7: grant lemmas 1,2,3 to the original problem and prove $(y^{2})^{x}$ divides $x^{x}$
  • Lemma 8: grant lemmas 1,2,3,4 to the original problem and prove $(y^{2})^{x}$ divides $x^{x}$
This latter process leads to new lemmas with proofs easier than the proof of the original problem. For example, the proof for lemma 5 would need to still prove lemmas 2,3,4 and then prove the ultimate goal. The proof for lemma 6 is shorter/easier than the proof for lemma 5 while it is longer than the proof for lemma 7, and so on.
Overall, these will lead to $2*k$ lemmas. Each of these lemmas may also be broken down if they have proofs that are longer than a certain threshold.
In an extreme case, a proof may have $k=(n-1)/3$ intermediate hypothesis (each of them with 1 line of statement and 2 lines of proof leading to the 3 in the denominator) which would lead to $2/3*(n-1)$ extracted lemmas. For example, in this setting, a proof with 4 intermediate hypothesis may have $4*3+1=13$ lines. And that leads to $2/3*(13-1)=8$ extracted lemmas, as explained above.
E.2 Case of a proof with no intermediate structure
Now, we look at an extreme case of a proof with $n$ lines that does not have any specific structure, and we also assume that the $n$ lines of proof do not have any intermediate hypotheses. In such a case, we perform a forward path and two rounds of backward paths on the proof.
The forward path would start with the first line of the proof, apply the line, obtain the new proof state, and define that as a new lemma. The proof for this new lemma is the $(n-1)$ lines of the proof that remain. Continuing this way, the forward path can extract $(n-2)$ lemmas where each lemma has at least 2 lines of proof.
The first backward path would take each pair of consecutive lines in the proof, take the changes in the state, and grant them back to the original state as a hypothesis. The proof for such lemmas would be the two aforementioned consecutive lines of proof. This can lead to $(n-2)$ additional lemmas.
The second backward path generates lemmas where their proofs are the first $m$ lines of the original $n$ line proof. This leads to $(n-3)$ additional lemmas.
Therefore, in this case, one can theoretically extract $(3*n-7)$ lemmas. For a proof with 3 lines, this leads to extraction of 2 distinct lemmas. For a proof with 4 lines, this leads to extraction of 5 distinct lemmas, and so on.
This is a somehow mechanistic view of the process of breaking down the proofs. In practice, we have used human judgment. On average, for all the IMO problems in the paper, we have extracted about $n/3.5$ lemmas for $n$ lines of proof which is far less than the two bounds explained above.
The smallest ratio, in our dataset, is $n/5.6$ for IMO 1962 P2, and the largest ratio is $n/2.4$ for IMO 2022 P2.
E.3 Statistics about our lemmas and their proofs
Table E2: Characteristics about the length of the proof of the lemmas in our dataset
# Problem Mean Max Min Std Lemma count Lines of Lean code for all lemmas
1 1959-P1 2.3 3.0 1.0 1.2 3 30
2 1960-P2 7.1 30.0 2.0 9.4 8 140
3 1962-P2 6.6 16.0 1.0 4.7 14 220
4 1964-P2 8.6 25.0 2.0 8.1 9 180
5 1965-P2 17.2 158.0 2.0 29.6 73 2,950
6 1983-P6 10.3 27.0 2.0 6.1 55 1,200
7 1984-P6 13.4 137.0 1.0 22.2 68 1,620
8 1992-P1 11.0 70.0 1.0 12.5 91 2,135
9 1997-P5 12.4 85.0 1.0 14.4 122 2,950
10 2022-P2 12.2 66.0 1.0 14.5 61 1,630
11 2022-P5 12.2 77.0 2.0 15.2 266 6,845
12 2023-P4 22.2 115.0 1.0 29.5 137 5,580
Total 907 25,480
Appendix F Datasheet
Following the framework in [12] and similar to [20], Table LABEL:tbl:datasheet provides the data sheet for our dataset.
Table F3: Datasheet for our dataset
Questions Answers
Motivation
For what purpose was the dataset created? To provide a set of formal mathematical problems as a stepping stone for automated systems to prove challenging International Math Olympiad problems in number theory and algebra.
Who created the dataset and on behalf of which entity? The authors of this work.
Who funded the creation of the dataset? The companies where the authors work.
Any other comment? None.
Composition
What do the instances that comprise the dataset represent? Lemmas formalized in Lean 4.
How many instances are there in total? 907 lemmas.
Does the dataset contain all possible instances or is it a sample of instances from a larger set? It is not a sample of a larger set.
What data does each instance consist of? A formal lemma and its proof formalized in Lean.
Is there a label or target associated with each instance? Not necessarily.
Is any information missing from individual instances? No.
Are relationships between individual instances made explicit? Not applicable. Each lemmas can be proved on its own, and it is not dependent or have an explicit relationship with other lemmas in the dataset.
Are there recommended data splits? The dataset does not have such splits. However, it is possible for people to generate such splits. The authors do not advocate for such splits on this particular dataset. Instead, it is recommended to solely rely on the lemmas in the Mathlib library to prove the lemmas in this dataset.
Are there any errors, sources of noise, or redundancies in the dataset? No. There are absolutely no errors in the dataset as all the lemmas and their proofs are automatically checked by the system of Lean theorem prover. Redundancies is hard to define in this context. There are no repeated lemmas in the dataset.
Is the dataset self-contained, or does it link to or otherwise rely on external resources? It is self contained. It is written in Lean 4 programming language and it relies on the Mathlib library which is open source.
Does the dataset contain data that might be considered confidential? No.
Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? No.
Collection Process
How was the data associated with each instance acquired? 12 IMO problems were chosen as described in Table 1. The formal proof for each of these problems was written manually by the authors. These proofs were broken into their building blocks leading to a set of 907 lemmas written in Lean 4. The proofs for these lemmas were written by the authors. Lemmas that can be proved by native automated solvers in Lean were excluded from the dataset. The said native solvers are the following: hint, linarith, exact?, simp, omega, ring, norm_cast and norm_num.
What mechanisms or procedures were used to collect the data? Dataset was created by the authors based on the mathematical proofs of 12 IMO problems listed in Table 1.
If the dataset is a sample from a larger set, what was the sampling strategy? Not applicable.
Who was involved in the data collection process and how were they compensated? Dataset was not collected.
Over what timeframe was the data collected? Dataset was not collected.
Were any ethical review processes conducted? No.
Preprocessing/cleaning/labeling
Was any preprocessing/cleaning/labeling of the data done? No. There was a post processing which excluded lemmas that could be proved directly via native Lean solvers. These lemmas were deemed easy to prove, and since existing solvers can prove them, we do not consider them an obstacle in our larger goal of proving IMO problems.
Was the “raw” data saved in addition to the preprocessed/cleaned/labeled data? Not applicable.
Is the software that was used to preprocess/clean/label the data available? Lean prover was used to verify the correctness of the process which is an open source system widely used in this field of research.
Any other comments? No.
Uses
Has the dataset been used for any tasks already? We have evaluated the ability of GPT-4 in proving the lemmas in our dataset.
Is there a repository that links to any or all papers or systems that use the dataset? There will be a link on GitHub, but we are not sharing it at this time.
What (other) tasks could the dataset be used for? The first goal would be to develop models and methods that can prove these lemmas in formal language and also in natural language. The second goal would be to develop models and methods that can use these lemmas to prove the IMO problems in this dataset as well as other IMO problems in number theory and algebra.
Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses? We do not foresee such issues arising.
Are there tasks for which the dataset should not be used? Not that we can think of.
Any other comments? No.
Distribution
Will the dataset be distributed to third parties outside of the entity on behalf of which the dataset was created? Yes, we plan to release and host the dataset on GitHub.
How will the dataset will be distributed? GitHub.
When will the dataset be distributed? With the camera-ready version of the paper.
Will the dataset be distributed under a copyright or other intellectual property license, and/or under applicable terms of use? Yes, we plan to release it under the MIT license.
Have any third parties imposed IP-based or other restrictions on the data associated with the instances? No.
Do any export controls or other regulatory restrictions apply to the dataset or to individual instances? No.
Any other comments? No.
Maintenance
Who will be supporting/hosting/maintaining the dataset? The first author.
How can the owner/curator/manager of the dataset be contacted? Email address and also via the available methods on GitHub.
Is there an erratum? This is not necessary. All the lemmas and their proofs are checked and verified for correctness with the Lean system.
Will the dataset be updated? Possibly. If a new version of Lean is released, e.g., Lean 5, we might need to update our dataset based on the new release of Lean and Mathlib.
If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances? Not applicable.
Will older versions of the dataset continue to be supported/hosted/maintained? Yes.
If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so? Yes. There are two avenues. First, the dataset written in Lean can be rewritten in other languages such as Isabelle. Second, more problems (IMO or otherwise) can be formalized in Lean and/or other languages, and their proofs can be turned into a set of additional lemmas to expand the current dataset.
Any other comments? No.
Appendix G Experimenting with GPT-4: Prompt structure and problem statements
Figure G1 shows an example of how we interact with GPT-4. For each of the lemmas, the prompt starts with the following statement:
Please write a formal proof for this problem in lean language. Please think step by step and explain all your reasoning
After this, the formal lean statement of the lemma is provided.
Figure G1: The structure of the prompt to GPT-4 followed by the formal lemma statement in Lean.
The responses of GPT-4 are graded both for correctness of its explanation in natural language and also for the correctness of its formal proof in Lean. The results were presented in Tables 23.
In cases where GPt-4’s answer is not correct, we follow up with a prompt explaining the mistakes that we detect in the answer, instructing GPT-4 to correct the mistakes. Figures G2 and G3 show examples of such interactions.
Figure G2: Example of hinting GPT-4 and giving feedback about its mistakes.
These interactions rarely lead to a correct answer from GPT-4 as mistakes tend to repeat even after they are explained to GPT-4, e.g., Figure G3. It appears that many of the mistakes and hallucination of lemmas are a byproduct of GPT-4 not having any frame of reference in its training set. Sometimes, after a few rounds of feedback, GPT-4 stops giving any answer and it explains that the problem is hard to prove. Table 4 presents the results of hinting GPT-4 up to 10 rounds.
For these cases, a sensible analogy might be a student who has not studied the materials and shows up at an exam. Such student might improvise and make up answers, answers that might be often nonsensical, drawing from nonexistent rules, principles, and theorems. Such student might also not provide any answers and explain that they are uneducated on the topics of the exam.
Figure G3: Example of hinting GPT-4 and giving feedback about its mistakes.
Appendix H Source Criticism
In the supplementary material, we show the results of using source criticism to compare the correctness of the GPT-4 responses with proofs available on the Internet. We report a strong statistical correlation between the availability of proofs on the Web and the correctness of GPT-4’s responses. To search the Web, we use the Google search engine.
Source criticism is a scientific method and research methodology used in many fields, including information sciences, empirical sciences, and other fields, especially those that work with text [13]. Since the advent of the Internet, a large amount of text has accumulated on the Web. Sometimes it is useful or necessary to trace a given piece of text back to its sources, such as in the case of fact-checking to overcome misinformation, or in legal cases related to copyright. When dealing with the large amounts of text on the Internet, source criticism has proven to be a useful method [27]. We believe this method can also be useful in evaluating the capabilities and generalization of LLMs.
Comparing current texts with previously created texts yields insight into the composition of the newer text. The comparison can tell whether a present text has copied from earlier text and if so how closely the copied text resembles the earlier text. The oldest known variant that informs the later texts qualifies as the primary source. Texts copied or derived from the primary source are secondary sources. A secondary source is not necessarily an exact copy of a primary source. Indeed, a secondary source/text may have novelty of its own by putting together text from two or more different primary sources, or by significantly altering a primary text.
In our case of theorem proving, the primary sources are relatively easy to locate online. For example, most of the publicly available formal proofs written in Lean language can be found on GitHub. LLMs such as ChatGPT [18] and Llama [28], and even specialized models like Llemma [4], ProofNet [3], and Code Llama [24] are trained on contents of the web: either directly trained on Common Crawl or fine tuned from a model that was previously trained on Common Crawl. Although the GPT-4 training set is kept as a secret, we know that previous iterations of this model, such as the GPT-f model, were trained on the Common Crawl and WebMath datasets [21]. The WebMath contains all publicly available code and text on GitHub, arXiv, and Math StackExchange. Hence, the content of the Web can be a reasonable proxy for their training set. Given this, source criticism presents a largely untapped opportunity to evaluate the generalization capabilities of LLMs.
When an LLM produces a proof for a given lemma merely by copying the proof from its training set, we can consider that as a successful direct retrieval. We do not consider such retrieval to be trivial or useless. In fact, correct retrieval of a proof can be a useful capability for a model.
In addition to direct retrieval from the training set, a model may also be able to draw from various sources in its training set and put together a correct response. This would be a more advanced capability compared to direct retrieval and we call that generalization. [^1] Sometimes, generalization is contrasted against memorization, in the machine learning literature and also in educational studies. Direct retrieval that we described here can be considered memorization. This might imply lack of understanding when it comes to learning. However, here, we prefer not to use the term memorization as it may create certain confusions in the context of machine learning, especially the way memorization is used in computer vision. To avoid confusions, we use the term direct retrieval. The source criticism method described above can distinguish between these cases. If a proof (or a piece of text) is a direct copy of a primary source, we can consider that a direct retrieval. However, when a proof for a given lemma does not exactly appear in a primary source, we consider that a more advanced form of behavior by the model and it would no longer be counted as direct retrieval. We consider this form of generalization in three forms:
  1. 1.
    The proof consists of various pieces that are taken from two or more primary sources. The model has taken these pieces from those sources and has managed to put them together in a sensible way.
  2. 2.
    The proof generated by the model does not exactly appear in any primary source, rather, it is an alteration of a proof from a primary source.
  3. 3.
    A combination of the above two forms. For example, a model has retrieved some segment of a proof from some primary source and has put it together with an alternation of some other proof segment.
Here, we only study the correct proofs written by GPT-4, and we would like to know what fraction of these correct proofs can be direct retrieval from its training set (considered to be the Web) and how many of those correct proofs are of the generalization form where the model has made some effort in generating the proof. This effort can be the change of the variable names to align the proof with the current lemma, or it can involve more complex modifications. The results of this analysis is presented in Table H4.
Table H4: Results of analyzing the sources of correct proofs written by GPT-4. An entry of "a/x" in the "direct retrieval" column implies that there were $x$ number of lemmas that GPT-4 were able to write a correct proof for them, and there were $a$ lemmas that can be considered a direct retrieval. Entries of -/0 refer to cases where GPT-4 were not able to write any correct proofs.
# Problem Direct retrieval Modifications from one primary source Complex generalization
1 1959-P1 -/0 -/0 -/0
2 1960-P2 0/1 1/1 0/1
3 1962-P2 1/5 3/5 1/5
4 1964-P2 0/1 1/1 0/1
5 1965-P2 2/8 5/8 1/8
6 1983-P6 -/0 -/0 -/0
7 1984-P6 -/0 -/0 -/0
8 1992-P1 -/0 -/0 -/0
9 1997-P5 -/0 -/0 -/0
10 2022-P2 -/0 -/0 -/0
11 2022-P5 -/0 -/0 -/0
12 2023-P4 -/0 -/0 -/0
Total count 3/15 10/15 2/15
Total percentage 20% 67% 13%
The training set of GPT-4 is not limited to the contents of the Internet. According to its technical report, its training set also contains human generated material, as well as contents obtained from user interactions with GPT-4 accumulated over time. Hence, when the model provides a correct proof and that proof is not found on the Internet, it may still be a direct retrieval from some part of the model’s training set that is not on the Web. This distinction is hard to detect and verify but when interpreting the results in Table H4, it would be useful to keep in mind that the cases of complex generalization might belong to other columns.
One may use source criticism to analyze the incorrect proofs as well. However, given the large amount of incorrect proofs that GPT-4 generated in our experiments, performing source criticism on them may be more labor intensive than insightful. Nevertheless, these are the failure modes of an AI model and studying these failure modes will be crucial when developing automated systems for theorem proving which falls outside the scope of the current paper.
Appendix I Proof repair
For proof repair, we take the incorrect proofs provided by GPT-4 and repair them manually. The goal is to see whether GPT-4’s proof can be useful starting points so that humans do not need to write the proof from scratch. To quantify the effort, we report different measures:
  1. 1.
    Percentage of lines that need to be edited
  2. 2.
    Whether the proof sketch by GPT-4 is correct
  3. 3.
    Levenshtein ratio between GPT-4’s proof and the correct proof
For item 1, if new lines have to be added, those will be counted towards the number of edited lines. As a result, this percentage can become larger than 100%, indicating that much is missing in the generated proof. The second measure aims to identify whether one can take the sketch of the proof from GPT-4 and fill in the structure. The third measure is easy to calculate and from it, we can see a correlation between this measure and the human effort needed to repair the proof (item 1). It can be useful for the community to look at this measure as a proxy for the effort of repairing proofs.
The results will be included in the supplementary material. In summary, for almost all of the lemmas, the effort of writing the proof from scratch is less than the effort needed to go through the GPT-4’s proof and fix the issues in it.